OpenStudy (anonymous):

dy/dx=2y^2 and if y=-1 when x=1, then when x=2, y=? thanks

OpenStudy (anonymous):

dy (y^-2) = dx -1/y = x + C y = -1 / (x + C) y(1) = -1 -1 / (1 + C) = -1 C = 0 y = -1 / x y = -1 / 2

OpenStudy (anonymous):

shoot its 2x

OpenStudy (anonymous):

-1/y = 2x + C y = -1 / (2x + C) y(1) = -1 -1 / (2 + C) = -1 C = -1 y = -1 / (x - 1) y = -1 / (2 - 1) y = -1

OpenStudy (anonymous):

isn't y=-1/3? dy/y^2=2 dx y^-2=2dx -1/y=2x+c c=-1 -1/y=2x-1 y=-1/3

OpenStudy (anonymous):

yea i found it to be -1/3

OpenStudy (anonymous):

THANKS ANYWAY

OpenStudy (anonymous):

Sorry, it is y = -1 / (2 - -1) = -1 / 3

OpenStudy (anonymous):

I forgot the second negative

OpenStudy (anonymous):

thanks